Please explain why the measure of a complement of an angle can never be exactly half the measure of the supplement of the angle
Answer by richard1234(7193)

What if the angle measures 0 degrees? Then the complement is 90 and the supplement is 180.

For angular measures greater than 0 and less than 90, this is impossible because if (180-x) = 2(90-x), then -x = -2x which doesn't occur unless x = 0.